# Stochastic Polyak Step-sizes and Momentum: Convergence Guarantees and Practical Performance

### v1

@@ Appendix D Proofs of the main results
Appendix E Additional Experiments
E.1 Extra Deterministic Experiment
Here we have included an extra experiment in the deterministic setting, this one for logistic regression. We have performed a grid search to find the best $\beta$ for MomPSmax since no optimal choices for $\beta$ are known for the general convex case. We use the same $\beta$ for HB and ALR-MAG for direct comparison. For the step-size of HB we used $\gamma=2(1-\beta)/L$ as recommended in Ghadimi et al. [2015]. The results are presented in Figure 7. In this problem we observe that our step-size (MomPSmax) is the fastest, having similar performace with ALR-MAG and GD Polyak. For the deterministic setting we have used a 12 core AMD Ryzen 5 5600H CPU to run the experiments.
@@ E.1 Extra Deterministic Experiment
E.2 Stochastic Heavy Ball with Constant Step-size
@@
Figure 8: Comparison of SHB with constant step-size. Liu et al. [2020] vs Corollary 3.4 for various momentum coefficients $\beta$ on logistic regression with synthetic data.
E.3 Other Choices
@@
E.4 Is $\gamma_{b}$ needed?
Recall the definition of MomPSmax, $\gamma_{t}=(1-\beta)\min\left\{\frac{f(x^{t})-f(x^{*})}{\|\nabla f(x^{t})\|^{2}},\gamma_{b}\right\}$. The more natural choice would be without the upper bound $\gamma_{b}$. Moreover, as in Section E.3, we can have the naive version of both of the above step-sizes, i.e. with no correcting factor $1-\beta$. We compare the following step-sizes on both the least squares problem, in Figure 12, and logistic regression problem, in Figure 13, in the deterministic setting on synthetic data. The method is SHB with $\beta=0.97$ for the least squares and $\beta=0.3$ for the logistic regression.
@@ E.5 Plots for $C_{1}$ and $C_{2}$
E.6 More convex experiments and parameter settings
@@
E.7 More deep learning experiments and parameter settings
In this section, we list the parameters, architectures and hardware that we used for the deep learning experiments. The information is collected in Table 3. We also include some extra experiments in Figures 19 and 20.
[…]
 Architecture ResNet 10/34 [He et al., 2016], Densenet 121 [Huang et al., 2017]
@@ E.7 More deep learning experiments and parameter settings
[…]
Table 3: CIFAR-10 experiment
Figure 19: Resnet 18 on CIFAR-10
Figure 20: DenseNet 121 on CIFAR-10

### v2

@@ Appendix D Proofs of the main results
Appendix E Beyond the bounded iterates assumption
The two main theorems on the convergence of MomDecSPS/MomAdaSPS require the bounded iterates assumption. As mentioned in the main paper, this is a standard assumption for several adaptive step-sizes, see: [Reddi et al., 2018, Ward et al., 2020, Orvieto et al., 2022, Jiang and Stich, 2023]. However, one can artificially remove this assumption by adding a projection step in the update rule of IMA onto a compact and convex subset $\mathcal{D}\subseteq\mathbb{R}^{d}$ as described in this section.
Consider the constrained finite-sum optimization problem,
$\min_{x\in\mathcal{D}}\left[f(x)=\frac{1}{n}\sum_{i=1}^{n}f_{i}(x)\right],$ (36)
where each $f_{i}:\mathbb{R}^{d}\to\mathbb{R}$ is convex, smooth, and lower bounded by $\ell_{i}^{*}$ and $\mathcal{D}\subseteq\mathbb{R}^{d}$ is a compact and convex subset. Let $X^{*}_{\mathcal{D}}\subseteq\mathcal{D}$ be the set of minimizers of (36). We assume that $X^{*}_{\mathcal{D}}\neq\emptyset$ and we fix $x^{*}\in X^{*}_{\mathcal{D}}$.
The new update rule of IMA, takes the following form (Projected IMA):
$z^{t+1}=\textnormal{proj}_{\mathcal{D}}\left[z^{t}-\eta_{t}\nabla f_{S_{t}}(x^{t})\right]$ (37)
with $x^{0}=z^{0}\in\mathcal{D}$, where $\textnormal{proj}_{\mathcal{D}}(x)\in\arg\min_{d\in\mathcal{D}}\|d-x\|^{2}$. Note that the projection step is only needed for the update of $z^{t+1}$ because $x^{t+1}$ is already a convex combination of elements in the convex set $\mathcal{D}$, namely $z^{t+1}$ and $x^{t}$ by induction. We highlight the fact that Projected IMA is not necessarily equivalent to Projected SHB. Now the proofs of Theorems D.2 and D.4 will go through using the non-expansiveness of the projection operator, as explained in the following lemma:
Lemma E.1 (Non-expansiveness).
For all $x,y\in\mathbb{R}^{d}$ it holds
$\|\textnormal{proj}_{\mathcal{D}}(x)-\textnormal{proj}_{\mathcal{D}}(y)\|\leq\|x-y\|$
Proof.
See Example 8.14 and Lemma 8.16 in Garrigos and Gower [2023]. ∎
Now we have:
Theorem E.2 (Pojected IMA version of Theorem D.2).
Assume that each $f_{i}$ is convex and $L_{i}$-smooth. Then, the iterates of Projected IMA with
$\operatorname{\mathbb{E}}[f(\overline{x}^{T})-f(x^{*})]\leq\frac{2\lambda_{1}[f(x^{0})-f(x^{*})]}{T}+\frac{c_{T-1}D^{2}}{T\alpha}+\frac{\sigma^{2}}{T}\sum_{t=0}^{T-1}\frac{1}{c_{t}},$
where $\overline{x}^{T}=\frac{1}{T}\sum_{t=0}^{T-1}x^{t}$, $\alpha=\min\left\{\frac{1}{2L},\eta_{b}c_{0}\right\}$ and $D=\textnormal{diam }\mathcal{D}$.
Proof.
We have
$\|z^{t+1}-x^{*}\|^{2}$ $=\|\textnormal{proj}_{\mathcal{D}}(z^{t}-\eta_{t}\nabla f_{S_{t}}(x^{t}))-\textnormal{proj}_{\mathcal{D}}(x^{*})\|$
${\leq}\|z^{t}-\eta_{t}\nabla f_{S_{t}}(x^{t})-x^{*}\|$
$=\|z^{t}-x^{*}\|^{2}-2\eta_{t}\langle\nabla f_{S_{t}}(x^{t}),z^{t}-x^{*}\rangle+\eta_{t}^{2}\|\nabla f_{S_{t}}(x^{t})\|^{2}.$
Now we continue exactly like the rest of the proof of Theorem D.2. ∎
Theorem E.3 (Pojected IMA version of Theorem D.4).
Assume that each $f_{i}$ is convex and $L_{i}$-smooth. Then, the iterates of Projected IMA with
where $\overline{x}^{T}=\frac{1}{T}\sum_{t=0}^{T-1}x^{t}$, $\tau=\left(\lambda_{1}\sqrt{f(x^{0})-f(x^{*})}+\frac{cLD^{2}}{2}+\frac{1}{2c}\right)$ and $D=\textnormal{diam }\mathcal{D}$.
Proof.
Same as above. ∎
Appendix F Additional Experiments
F.1 Extra Deterministic Experiment
Here we have included an extra experiment in the deterministic setting, this one for logistic regression. We have performed a grid search to find the best $\beta$ for MomPSmax since no optimal choices for $\beta$ are known for the general convex case. We use the same $\beta$ for HB and ALR-MAG for direct comparison. For the step-size of HB we used $\gamma=2(1-\beta)/L$ as recomended in Ghadimi et al. [2015]. The results are presented in Figure 7. In this problem we observe that our step-size (MomPSmax) is the fastest, having similar performace with ALR-MAG and GD Polyak. For the deterministic setting we have used a 12 core AMD Ryzen 5 5600H CPU to run the experiments.
@@ F.1 Extra Deterministic Experiment
F.2 Stochastic Heavy Ball with Constant Step-size
@@
Figure 8: Comparison of SHB with constant step-size. [Liu et al., 2020] vs Corollary 3.5 for various momentum coefficients $\beta$ on logistic regression with synthetic data.
F.3 Other Choices
@@
F.4 Is $\gamma_{b}$ needed?
Recall the definition of MomPSmax, $\gamma_{t}=(1-\beta)\min\left\{\frac{f(x^{t})-f(x^{*})}{\|\nabla f(x^{t})\|^{2}},\gamma_{b}\right\}$. The more natural choice would be without the upper bound $\gamma_{b}$. Moreover, as in Section F.3, we can have the naive version of both of the above step-sizes, i.e. with no correcting factor $1-\beta$. We compare the following step-sizes on both the least squares problem, in Figure 12, and logistic regression problem, in Figure 13, in the deterministic setting on synthetic data. The method is SHB with $\beta=0.97$ for the least squares and $\beta=0.3$ for the logistic regression.
@@ F.5 Plots for $C_{1}$ and $C_{2}$
F.6 More convex experiments and parameter settings
@@
F.7 More deep learning experiments and parameter settings
In this section, we list the parameters, architectures and hardware that we used for the deep learning experiments. The information is collected in Table 3. We also include some extra experiments in Figures 19 and 20. For ALR-SMAG we use $c=0.1$ in the DL experiments.
Hyper-parameter Value
Architecture ResNet 10/34 [He et al., 2016]
@@ F.7 More deep learning experiments and parameter settings
Table 3: CIFAR10 experiment
Figure 19: Resnet 18 on CIFAR 10
Figure 20: Resnet 18 on CIFAR 100